.0180 +/- .0Z0


1
Expert's answer
2021-10-25T16:28:06-0400

1.

"0.0180-0.020=-0.002"

"0.0180+0.020=0.038"

2.


"0.0270-0.010=0.017"

"0.0270+0.010=0.037"